a man whose mass is 100kg climbs a tree in 20 seconds and expends a power at 150 watts. how tall is the tree
1
Expert's answer
2019-06-11T10:11:06-0400

The power

"P=\frac{\rm{work\; done}}{\rm{time}}=\frac{mgh}{t}"

So, the height of a tree

"h=\frac{Pt}{mg}""=\frac{150\:\rm{W}\times 20\:\rm{s}}{100\:\rm{kg}\times 10\:\rm{N/kg}}=3\:\rm{m}"